# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ Abstract
We argue that progress toward general intelligence requires complementary foundation models grounded in language, the physical world, and structured data. This report presents LimiX, the first installment of our large structured-data models (LDMs). LimiX treats structured data as a joint distribution over variables and missingness, thus capable of addressing a wide range of tabular tasks through query-based conditional prediction via a single model. LimiX is pretrained using masked joint-distribution modeling with an episodic, context-conditional objective, where the model predicts for query subsets conditioned on dataset-specific contexts, supporting rapid, training-free adaptation at inference. We evaluate LimiX across 10 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, class number, categorical–to-numerical feature ratio, missingness and sample-to-feature ratios. With a single model and a unified interface, LimiX consistently surpasses strong baselines including gradient-boosting trees, deep tabular networks, recent tabular foundation models, and automated ensembles, as shown in Figure 1 and Figure 2. The superiority holds across a wide range of tasks, such as classification, regression, missing value imputation, and data generation, often by substantial margins, while avoiding task-specific architectures or bespoke training per task. All LimiX models are publicly accessible under Apache 2.0.
@@ 1 Introduction
We posit that progress toward general intelligence is best organized around three complementary spaces: language, physical-world, and structured data, each anchored to a distinct data modality and set of inductive biases. In the language space, large language models (LLMs) provide a universal interface for natural and programming languages and have rapidly advanced instruction following, tool use, and explicit reasoning over token sequences (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Bai et al., 2023). In the physical-world space, recent foundation models ground knowledge in space perception and embodied reasoning. They emphasize spatial intelligence through structured scene understanding, controllable scene generation, and neural radiance field reconstruction (Mildenhall et al., 2020; Ke et al., 2025; Xiang et al., 2025; Li et al., 2024a; Yi et al., 2024). These models also include self-supervised video world models such as V-JEPA, which learn predictive abstractions from large-scale videos and support downstream planning and control (Bardes et al., 2024; Assran et al., 2025).
@@
On the other hand, structured data serves as the foundational bedrock for evidence-based decision-making across a multitude of critical domains, including finance, healthcare, logistics, and public policy (Fuster et al., 2022; Johnson et al., 2016; Yu et al., 2021; Krafft et al., 2020). The structural consistency and inherent order of structured data provide a robust framework for quantitative analysis and reliable operations (Ramakrishnan et al., 2003; Silberschatz et al., 2011; Stonebraker & Çetintemel, 2018), enabling precise prediction, automated reasoning, and rigorous causal inference (Pearl, 2009; Hernán & Robins, 2010; Little & Rubin, 2019). While the emergence of unstructured data has captured considerable attention, the analytical power and operational reliability of structured data remain unparalleled for a vast array of real-world applications (Fang et al., 2024; Van Breugel & Van Der Schaar, 2024). Consequently, advancements in structured-data prediction, analysis, and reasoning are not merely an academic pursuit but a critical enabler for efficiency, innovation, and accuracy in modern data-driven systems. Structured data is not subsumed by language models or embodied intelligence. Converting tables into free text discards metric geometry, physical units, and patterns of missingness that are central to reliable prediction, while models focusing on perception and control in three-dimensional physical environments do not capture discrete interventions, business rules, or causal heterogeneity across environments. Empirical surveys also document current limitations of language models on tabular prediction without bespoke adaptation (Fang et al., 2024; Sui et al., 2024).
@@
Traditionally, practitioners deploy pipelines of specialized models with gradient-boosting trees and automated ensembles such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), CatBoost (Dorogush et al., 2018), and AutoGluon (Erickson et al., 2020)—that are trained separately for each dataset and task. These systems excel at supervised prediction but require full retraining on every new dataset, prolonging deployment and preventing reuse of knowledge across domains. Recent deep approaches for tables have improved accuracy on mixed-type data (Huang et al., 2020; Yoon et al., 2020; Gorishniy et al., 2021; Somepalli et al., 2022; Bahri et al., 2022). However, they are still typically trained per dataset and do not provide a single model that transfers across objectives and constraints.
@@
These limitations motivate the pursuit of a foundation model for structured data, i.e. models trained on large families of datasets to perform in-context learning (ICL) without per-task fine-tuning. Notably, TabPFN (Hollmann et al., 2022) and its successor TabPFN-v2 (Hollmann et al., 2025) demonstrate state-of-the-art performance and speed on small-to-medium-scale tables via prior-data fitting over diverse generative processes, and community efforts increasingly argue for tabular foundation models as a distinct paradigm (Van Breugel & Van Der Schaar, 2024). Concurrent works such as TabDPT (Ma et al., 2024) and TabICL (Qu et al., 2025) explore scaling these ideas to real and larger datasets, narrowing the gap between tabular foundation models and classical methods under the scenarios of larger sample sizes.
@@
In this work, we introduce LimiX, the first installment of our LDM series. LimiX aims to push generality further: a single model capable of classification, regression, missing value imputation, data generation, and sample selection for interpretability under one training and inference recipe, shifting the paradigm from bespoke pipelines to unified, foundation-style tabular learning. LimiX adopts a lightweight, scalable architecture that represents structured data as a set of sample–feature embeddings and learns dependencies across two dimensions: across features (columns) and across samples (rows). To make the attention module explicitly column-aware without inflating parameters, we introduce a low-rank discriminative feature encoding that encodes feature identities. Pretraining of LimiX follows a masked joint-distribution objective and an episodic, context-conditional formulation: For each dataset, an in-context subset establishes dataset-specific priors, and the model is trained to predict masked entries in a disjoint query subset, enabling per-dataset adaptation without fine-tuning at inference. The pretraining corpus consists of data synthesized from hierarchical structural causal models (SCMs). Within the synthesis pipeline, we employ graph-aware sampling to obey the causal structure and solvability-aware sampling to accommodate the data quantity of various downstream tasks, improving coverage of local patterns and generalization. At inference, attention-guided retrieval provides an efficient, optional ensemble and fine-tuning mechanism. LimiX retrieves informative samples and features using its own attention scores, aggregates predictions across a handful of lightweight pipelines, and delivers calibrated outputs for various downstream tasks, all through a unified conditional-inference interface and without task-specific architectures or bespoke per-dataset training.
@@
We evaluate LimiX on 10 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, number of classes, categorical–to-numerical feature ratios, missing values, and sample-to-feature ratios. Results show that LimiX demonstrates the strongest predictive performance on all the benchmarks. With a single model and a unified inference interface, LimiX surpasses competitive baselines including gradient-boosting trees, deep tabular networks, recent tabular foundation models, and automated ensemble methods. Across classification, regression, missing value imputation, data generation, and out-of-distribution prediction, LimiX delivers consistent gains, often by large margins, while avoiding task-specific model architectures, customized ensembles, or per-dataset training. On most benchmarks, such as OpenML-CC18 (Bischl et al., 2017), TabArena (Erickson et al., 2025), TALENT-REG (Liu et al., 2024), LimiX is the only model that consistently outperforms AutoGluon, which is considered an outstanding baseline across various tabular-data tasks.
@@ 2 Architecture
We consider a dataset $\mathcal{D}=\{({\mathbf{x}}^{R},{\mathbf{y}}^{R})\}$ of $d$ features and an outcome variable, where ${\mathbf{x}}^{R}=\{{\mathbf{x}}^{R}_{i}\}_{i=1}^{m}$, ${\mathbf{y}}^{R}=\{y^{R}_{i}\}_{i=1}^{m}$ and the superscript $R$ indicates the raw input. Here, ${\mathbf{x}}^{R}_{i}\in\mathbb{R}^{d}$ and $y^{R}_{i}\in\mathbb{R}$ correspond to the $i^{th}$ sample, while ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ correspond to the 2D tabular data. For in-context samples and test samples, we use subscripts to distinguish between them. For example, ${\mathbf{x}}^{R}_{ct}\in\mathbb{R}^{m_{ct}\times d}$ denotes the features of in-context samples, while ${\mathbf{x}}^{R}_{te}\in\mathbb{R}^{m_{te}\times d}$ denotes those of the test samples.
@@ 2.1 Embedding of Tabular Data
To ensure compatibility with modern architectures such as Transformers (Vaswani et al., 2017), each cell of the 2D tabular input $x^{R}_{i,j}\in\mathbb{R}$ is first projected into a latent embedding space $\mathbb{R}^{p}$. Specifically, ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ is transformed into ${\mathbf{x}}\in\mathbb{R}^{m\times d\times p}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ is transformed into ${\mathbf{y}}\in\mathbb{R}^{m\times p}$. All subsequent attention operations are then conducted within this latent embedding space. Such a high-dimensional embedding space could strengthen the expressivity of the model. Concretely, we employ a two-layer MLP with LayerNorm (Ba et al., 2016) as the embedding module, i.e. ${\mathbf{x}}_{i,j}=\sigma(LN(\sigma(LN({\mathbf{x}}^{R}_{i,j}{\mathbf{W}}^{(1)}+{\mathbf{b}}^{(1)})){\mathbf{W}}^{(2)}+{\mathbf{b}}^{(2)}))$, where LN is LayerNorm and $\sigma$ is the GELU activation function (Hendrycks & Gimpel, 2016). Separate embedding modules are used for ${\mathbf{x}}^{R}$ and ${\mathbf{y}}^{R}$.
Figure 3: The overall model structure of LimiX.
@@ 2.2 Discriminative Feature Encoding
The attention score between feature $j$ and $j^{\prime}$ for sample $i$ is $\frac{1}{\sqrt{p}}\left\langle{\mathbf{x}}_{i,j}{\mathbf{W}}_{\mathrm{Q}},{\mathbf{x}}_{i,j^{\prime}}{\mathbf{W}}_{\mathrm{K}}\right\rangle$. This score depends solely on the interactions between the embeddings of cell values and imposes no explicit prior on feature (column) identity. As a result, the model cannot reliably infer the column from which a cell value originates. Inspired by Hollmann et al. (2025), we introduce a learnable low-rank column identifier termed discriminative feature encoding (DFE) whose design philosophy includes two objectives: (i) Encodings of different features should be well separated to ensure discriminability; (ii) Encodings in the embedding space admit a low effective rank so that the model can express column identities compactly and share statistical strength across features.
@@ 2.3 Model Structure
As shown in Figure 3, LimiX comprises 12 transformer blocks. Each block performs axis-wise self-attention along the feature axis twice and along the sample axis once, and each attention sublayer is followed by a position-wise feed-forward network (FFN). We adopt this asymmetric design with two feature-level passes and one sample-level pass because ablations show that using equal numbers underrepresent feature interactions, whereas doubling attention along the feature axis improves modeling capacity on heterogeneous schemas with minimal overhead. All sublayers use LayerNorm (Ba et al., 2016) in a pre-normalization configuration to stabilize optimization and support scaling in depth.
@@ 3.1 Context-Conditional Masked Modeling for Joint Distribution Learning
We pretrain LimiX by randomly masking cells in each row and training the model to recover the hidden entries from the visible context. Exposing the model to various mask patterns forces it to master a wide spectrum of conditional dependencies among variables. When these conditionals are learned properly, they effectively define a single joint model of the data. This joint model can then be queried for diverse tasks with one mechanism: Treat a chosen column as the target to perform regression or classification, fill in missing values by predicting the masked cells from the observed ones, and generate new samples by iteratively masking and refilling subsets of features.
@@
Unlike BERT-style masked modeling (Devlin et al., 2019), where adaptation is implicit in parameters accumulated during pretraining, CCMM considers context as a non-parametric memory that the model can consult during inference. This yields per-dataset calibration, better handling of rare categories, and improved robustness to distribution shift without gradient updates, aligning with principles of in-context learning and meta-learning for mixed-type tabular data.
We also find that CCMM improves modeling of the underlying dependency structure compared with recent tabular foundation models (Hollmann et al., 2025; Qu et al., 2025). By demanding consistency across numerous conditional predictions of all features rather than optimizing the loss of a single conditional prediction of a prefixed feature, the model is compelled to capture stable variable–variable relations instead of brittle decision boundaries. As the coverage of masks becomes richer, these cross-conditional constraints get tightened, yielding more reliable recovery of the joint distribution of all the features and more stable estimates in downstream usages. Please refer to Section 6 for details.

### v2

@@ Abstract
We argue that progress toward general intelligence requires complementary foundation models grounded in language, the physical world, and structured data. This report presents LimiX-16M and LimiX-2M, two instantiations of our large structured-data models (LDMs). Both models treat structured data as a joint distribution over variables and missingness, thus capable of addressing a wide range of tabular tasks through query-based conditional prediction via a single model. They are pretrained using masked joint-distribution modeling with an episodic, context-conditional objective, supporting rapid, training-free adaptation at inference. We evaluate LimiX models across 11 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, class number, categorical–to-numerical feature ratio, missingness and sample-to-feature ratios. LimiX-16M consistently surpasses strong baselines, as shown in Figure 1 and Figure 2. The superiority holds across a wide range of tasks, such as classification, regression, missing value imputation, and data generation, often by substantial margins, while avoiding task-specific architectures or bespoke training per task. Notably, LimiX-2M delivers strong results under tight compute and memory budgets. We also present the first scaling law study for LDMs, revealing how data and model scaling jointly influence downstream performance and offering quantitative guidance for tabular foundation modeling. All LimiX models are publicly accessible under Apache 2.0.
@@ 1 Introduction
We posit that progress toward general intelligence is best organized around three complementary spaces: language, physical-world, and structured data, each anchored to a distinct data modality and set of inductive biases. In the language space, large language models (LLMs) provide a universal interface for natural and programming languages and have rapidly advanced instruction following, tool use, and explicit reasoning over token sequences (openai2023gpt4; touvron2023llama; team2023gemini; bai2023qwen). In the physical-world space, recent foundation models ground knowledge in space perception and embodied reasoning. They emphasize spatial intelligence through structured scene understanding, controllable scene generation, and neural radiance field reconstruction (mildenhall2020nerf; ke20243d; xiang2025repurposing; li2024behavior; yi2024gaussiandreamer). These models also include self-supervised video world models such as V-JEPA, which learn predictive abstractions from large-scale videos and support downstream planning and control (bardes2024vjepa; assran2025vjepa2).
@@
On the other hand, structured data serves as the foundational bedrock for evidence-based decision-making across a multitude of critical domains, including finance, healthcare, logistics, and public policy (fuster2022predictably; johnson2016mimic; yu2021scf; krafft2020defining). The structural consistency and inherent order of structured data provide a robust framework for quantitative analysis and reliable operations (ramakrishnan2003dbms; silberschatz2011db; stonebraker2018one), enabling precise prediction, automated reasoning, and rigorous causal inference (pearl2009causality; hernan2010causal; little2019statistical). While the emergence of unstructured data has captured considerable attention, the analytical power and operational reliability of structured data remain unparalleled for a vast array of real-world applications (fang2024llmtabular; vanbreugel2024priority). Consequently, advancements in structured-data prediction, analysis, and reasoning are not merely an academic pursuit but a critical enabler for efficiency, innovation, and accuracy in modern data-driven systems. Structured data is not subsumed by language models or embodied intelligence. Converting tables into free text discards metric geometry, physical units, and patterns of missingness that are central to reliable prediction, while models focusing on perception and control in three-dimensional physical environments do not capture discrete interventions, business rules, or causal heterogeneity across environments. Empirical surveys also document current limitations of language models on tabular prediction without bespoke adaptation (fang2024llmtabular; sui2024table).
@@
Traditionally, practitioners deploy pipelines of specialized models with gradient-boosting trees and automated ensembles such as XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), CatBoost (dorogush2018catboost), and AutoGluon (erickson2020autogluon)—that are trained separately for each dataset and task. These systems excel at supervised prediction but require full retraining on every new dataset, prolonging deployment and preventing reuse of knowledge across domains. Recent deep approaches for tables have improved accuracy on mixed-type data (huang2020tabtransformer; yoon2020vime; gorishniy2021revisiting; somepalli2021saint; bahri2021scarf). However, they are still typically trained per dataset and do not provide a single model that transfers across objectives and constraints.
@@
These limitations motivate the pursuit of a foundation model for structured data, i.e. models trained on large families of datasets to perform in-context learning (ICL) without per-task fine-tuning. Notably, TabPFN (hollmann2022tabpfn) and its successor TabPFN-v2 (hollmann2025tabpfn_nature) demonstrate state-of-the-art performance and speed on small-to-medium-scale tables via prior-data fitting over diverse generative processes, and community efforts increasingly argue for tabular foundation models as a distinct paradigm (vanbreugel2024priority). Concurrent works such as TabDPT (ma2024tabdpt) and TabICL (qu2025tabicl) explore scaling these ideas to real and larger datasets, narrowing the gap between tabular foundation models and classical methods under the scenarios of larger sample sizes.
@@
In this work, we present LimiX, a unified family of our LDM series, and release its first two variants, LimiX-16M and LimiX-2M. LimiX models aim to push generality further: a single model capable of classification, regression, missing value imputation, data generation, and sample selection for interpretability under one training and inference recipe, shifting the paradigm from bespoke pipelines to unified, foundation-style tabular learning. LimiX models adopt a lightweight, scalable architecture that represents structured data as a set of sample–feature embeddings and learns dependencies across two dimensions: across features (columns) and across samples (rows). To make the attention module explicitly column-aware without inflating parameters, we introduce a low-rank discriminative feature encoding that encodes feature identities. Pretraining of LimiX models follows a masked joint-distribution objective and an episodic, context-conditional formulation: For each dataset, an in-context subset establishes dataset-specific priors, and the model is trained to predict masked entries in a disjoint query subset, enabling per-dataset adaptation without fine-tuning at inference. The pretraining corpus consists of data synthesized from hierarchical structural causal models (SCMs). Within the synthesis pipeline, we employ graph-aware sampling to obey the causal structure and solvability-aware sampling to accommodate the data quantity of various downstream tasks, improving coverage of local patterns and generalization. At inference, attention-guided retrieval provides an efficient, optional ensemble and fine-tuning mechanism. LimiX models retrieve informative samples and features using their own attention scores, aggregate predictions across a handful of lightweight pipelines, and deliver calibrated outputs for various downstream tasks, all through a unified conditional-inference interface and without task-specific architectures or bespoke per-dataset training.
@@
We evaluate LimiX-16M and LimiX-2M across 11 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, class number, categorical–to-numerical feature ratio, missingness and sample-to-feature ratios. With a single model and a unified inference interface, LimiX-16M surpasses competitive baselines including gradient-boosting trees, deep tabular networks, recent tabular foundation models, and automated ensemble methods. Across classification, regression, missing value imputation, data generation, and out-of-distribution prediction, LimiX-16M delivers consistent gains, often by large margins, while avoiding task-specific model architectures, customized ensembles, or per-dataset training. Notably, LimiX-2M deliver strong results under tight compute and memory budgets.
On most benchmarks, such as OpenML-CC18 (bischl2017openml), TabArena (erickson2025tabarena), TALENT (liu2024talenttabularanalyticslearning), LimiX-16M is the only model that consistently outperforms AutoGluon, which is considered an outstanding baseline across various tabular-data tasks.
This work introduces the first scaling law for LDMs, providing quantitative insights into how data and model scaling shape downstream performance and inform principled design of tabular foundation models.
@@ 2 Architecture
We consider a dataset $\mathcal{D}=\{({\mathbf{x}}^{R},{\mathbf{y}}^{R})\}$ of $d$ features and an outcome variable, where ${\mathbf{x}}^{R}=\{{\mathbf{x}}^{R}_{i}\}_{i=1}^{m}$ and ${\mathbf{y}}^{R}=\{y^{R}_{i}\}_{i=1}^{m}$; the superscript $R$ indicates the raw input. Here, ${\mathbf{x}}^{R}_{i}\in\mathbb{R}^{d}$ and $y^{R}_{i}\in\mathbb{R}$ correspond to the $i^{th}$ sample, while ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ correspond to the 2D tabular data. For in-context samples and test samples, we use subscripts to distinguish between them. For example, ${\mathbf{x}}^{R}_{ct}\in\mathbb{R}^{m_{ct}\times d}$ denotes the features of in-context samples, while ${\mathbf{x}}^{R}_{te}\in\mathbb{R}^{m_{te}\times d}$ denotes those of the test samples.
@@ 2.1 Embedding of Tabular Data
To ensure compatibility with modern architectures such as Transformers (vaswani2017attention), each cell of the 2D tabular input $x^{R}_{i,j}\in\mathbb{R}$ is first projected into a latent embedding space $\mathbb{R}^{p}$. Specifically, ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ is transformed into ${\mathbf{x}}\in\mathbb{R}^{m\times d\times p}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ is transformed into ${\mathbf{y}}\in\mathbb{R}^{m\times p}$. All subsequent attention operations are then conducted within this latent embedding space. Such a high-dimensional embedding space could strengthen the expressivity of the model. Concretely, we employ a two-layer MLP with LayerNorm (ba2016layernorm) as the embedding module, i.e. ${\mathbf{x}}_{i,j}=\sigma(LN(\sigma(LN({\mathbf{x}}^{R}_{i,j}{\mathbf{W}}^{(1)}+{\mathbf{b}}^{(1)})){\mathbf{W}}^{(2)}+{\mathbf{b}}^{(2)}))$, where LN is LayerNorm and $\sigma$ is the GELU activation function (hendrycks2016gaussian). Separate embedding modules are used for ${\mathbf{x}}^{R}$ and ${\mathbf{y}}^{R}$.
Figure 3: The overall model structure of LimiX-16M.
@@ 2.2 Discriminative Feature Encoding
The attention score between feature $j$ and $j^{\prime}$ for sample $i$ is $\frac{1}{\sqrt{p}}\left\langle{\mathbf{x}}_{i,j}{\mathbf{W}}_{\mathrm{Q}},{\mathbf{x}}_{i,j^{\prime}}{\mathbf{W}}_{\mathrm{K}}\right\rangle$. This score depends solely on the interactions between the embeddings of cell values and imposes no explicit prior on feature (column) identity. As a result, the model cannot reliably infer the column from which a cell value originates. Inspired by hollmann2025tabpfn_nature, we introduce a learnable low-rank column identifier termed discriminative feature encoding (DFE) whose design philosophy includes two objectives: (i) Encodings of different features should be well separated to ensure discriminability; (ii) Encodings in the embedding space admit a low effective rank so that the model can express column identities compactly and share statistical strength across features.
@@ 2.3 Model Structure
As shown in Figure 3, LimiX models comprise 12 transformer blocks. Each block performs axis-wise self-attention along the feature and sample axes, incorporating position-wise feed-forward networks (FFNs). In LimiX-16M, we employ an asymmetric configuration with two feature-level passes and one sample-level pass, each followed by an FFN, as ablation studies demonstrate that equal numbers of feature and sample passes underrepresent feature interactions, while increased feature-axis attention enhances modeling capacity for heterogeneous schemas with negligible overhead. All sublayers use pre-normalized LayerNorm (ba2016layernorm) to stabilize optimization and support deeper scaling.
@@ 3.1 Context-Conditional Masked Modeling for Joint Distribution Learning
We pretrain LimiX models by randomly masking cells in each row and training the model to recover the hidden entries from the visible context. Exposing the model to various mask patterns forces it to master a wide spectrum of conditional dependencies among variables. When these conditionals are learned properly, they effectively define a single joint model of the data. This joint model can then be queried for diverse tasks with one mechanism: Treat a chosen column as the target to perform regression or classification, fill in missing values by predicting the masked cells from the observed ones, and generate new samples by iteratively masking and refilling subsets of features.
@@
Unlike BERT-style masked modeling (devlin2019bert), where adaptation is implicit in parameters accumulated during pretraining, CCMM considers context as a non-parametric memory that the model can consult during inference. This yields per-dataset calibration, better handling of rare categories, and improved robustness to distribution shift without gradient updates, aligning with principles of in-context learning and meta-learning for mixed-type tabular data.
We also find that CCMM improves modeling of the underlying dependency structure compared with recent tabular foundation models (hollmann2025tabpfn_nature; qu2025tabicl). By demanding consistency across numerous conditional predictions of all features rather than optimizing the loss of a single conditional prediction of a prefixed feature, the model is compelled to capture stable variable–variable relations instead of brittle decision boundaries. As the coverage of masks becomes richer, these cross-conditional constraints get tightened, yielding more reliable recovery of the joint distribution of all the features and more stable estimates in downstream usages. Please refer to Section 6 for details.